OpenStudy (anonymous):

Convert the polar equation to rectangular coordinates: r=8sec(theta)

OpenStudy (anonymous):

mutiply both side by r^2 =r*8sec(theta) but r^2= hence x^2+y^2=8rsec(theta) hence x^2+y^2-8rsec(theta)=0

OpenStudy (anonymous):

how would use that to find the rectangular coordinates?

OpenStudy (anonymous):

i managed to find x which is 8 by substituting rcos(theta) into the original

OpenStudy (anonymous):

but i'm struggling to find y

OpenStudy (anonymous):

put your answer of x in my answer you will get answer for y..

OpenStudy (anonymous):

but you have three variables: x, y, and theta. i only have x